Solve the equation by first clearing the fractions or decimals.
49.1/2t + 5/12 =5/6
Any help with this equation would be Greatly appreciated.

Answer by stanbon(75887)

1/2t + 5/12 =5/6
Multiply both sides by 12 to get:
6t + 5 = 2*5
6t = 5
t = 5/6
==============
Cheers,
Stan H.